Circular No. 6425

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

CM DRACONIS E. L. Martin and H. Deeg, Instituto de Astrofisica de Canarias, re-

port: “The TEP (Transits of Extrasolar Planets) international collaboration (working group: M. Chevreton, H. Deeg, L. Doyle, J. Jenkins, W. Lee, E. L. Martin, E. Palaiologou, J. Schneider) has been monitoring CM Dra since the first half of 1994 using CCD cameras on eight telescopes, with the aim of detecting planetary transits in this eclipsing-binary system. So far, ∼ 500 hr of effective integration time have been analyzed in some 16 000 CCD frames. Night-to-night variations have been usually < 0.02 mag in the R band. In an 8.2-hr observation in photometric conditions with the IAC80 telescope at Teide Observatory beginning 1994 May 27, the R brightness of CM Dra started at 0.08 mag below the normal zero level and gradually increased to normal in 4 hr; in the next two hours it decreased again by 0.04 mag, and in the last two hours it returned to its normal level (Deeg et al. 1996, A.Ap. Trans., in press). The first light drop is of amplitude similar to that reported on IAUC 6423 and could be due to the same planet; the period would be ∼ 735 days or a submultiple of it. However, caution must be exercised because confirmation of a planetary occultation requires repeated observations of transits in at least two different filters and the determination of the total duration and shape, which depend in a predictable way on the phase of the binary orbit (Jenkins et al. 1996, Icarus 119, 244). We also note that timing of CM Dra eclipses made by the TEP group in 1994–1996 shows that there is a mean delay of ∼ 130 s with respect to the ephemerides of Lacy (1977, Ap.J. 218, 444), though no timing periodicity is apparent in our data. We encourage interested observers to contact one of the TEP members for coordination of observations.”

SUPERNOVA 1996ai IN NGC 5005 C. Bottari, Sava, Italy, reports the following estimated magnitudes of

SN 1996ai from his CCD images (obtained with an HPC1 camera with a TC 215 chip and an ISIS Deep-Sky filter, which transmits at Hβ 486-nm, O III 501-nm, and Hα 656-nm): 1995 May 5.88, not visible; 1996 June 16.88, 14.5; 20.91, 13.8.

Corrigendum. On IAUC 6422, line 19, for V = 16.2 ± 0.2.” read V = 15.2 ± 0.2.”

1996 June 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green